OpenStudy (anonymous):

Compute the sum: S=1/2!+2/3!+ ....+n/(n+1)!

OpenStudy (jamesj):

Calculate a few partial sums S1 = 1/2! = 1/2 S2 = 1/2! + 2/3! = 1/2 + 1/3 = 5/6 S3 = ... = 23/24 S4 = ... = 119/120 See the pattern?